The real risk-free rate is 4%. Inflation is expected to be 3% this year, 4% next year, and then 3% thereafter. The maturity risk premium is estimated to be 0.0003 x (t - 1), where t = number of years to maturity. What is the nominal interest rate on a 7-year Treasury security? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Solution :-

We can use the following formula to compute the yield on 7-year treasury note:

Treasury note yield = Real Risk Free Rate + Average Rate of Inflation + Maturity Risk Premium

In this question given ,

The Real Risk Free Rate = 4%,

The Average Rate of Inflation for the next 7 years = [ 3% + 4% + (3% * 5) ] / 7 = 22% / 7 = 3.143%.

The Maturity Risk Premium = 0.0003 * ( t - 1 ) =  0.0003*(7 - 1) = 0.0018 = 0.18%.

Applying the formula, the yield on treasury note is:

Treasury note yield = 4% + 3.143% + 0.18%

Treasury note yield = 7.32%

If there is any doubt please ask in comments

Thank you Please Rate .............